|
Run summary
Shift logbook
Run settings
Beam logbook
BPD logbook
DAQ logbook
DCS logbook
Experts logbook
Gas logbook
TOF logbook
Trigger logbook
PSD logbook
VD logbook
Run coordination
Detector Upgrade
DRS logbook
|
| Logbook for beam experts |
 |
|
|
Message ID: 263
Entry time: Wed, 26.07.2023 22:41:12
|
| Author: |
Y. Balkova |
| Subject: |
Beam tuning K@60 GeV |
|
|
Wed, 26.07.2023 22:41:27 YB: After the magnets were ramped back up and MD was finished, I began to tune the beam around 19:00. The beam alignment was good. I put THC pressure to the theoretical kaon value. However, I had problems with CEDAR alignment as PMT1 was not working and PMT2 was used as trigger. Inaki intervened after Nikos created a ticket about the problem. He reset the default values for the delay between PMTs and the scintillator, and the issue was fixed. By the time he did that, I managed to align CEDAR with 6 PMTs. First I began the full scan (att. 1) to confirm that we see smth. Around seeing kaon peak, I terminated this scan and began another one to separate kaons from pions. During the scan, the beam moved, so I needed to terminate this scan and align CEDAR again.
pi-K scan:
PRESS TRIGG 6-FOLD 7-FOLD 8-FOLD
------------------------------------------------------------------------
1.8488 1.6669e+06 3.9643e-03 1.5178e-04 1.5598e-05
1.8292 1.6667e+06 2.1093e-02 3.8410e-03 7.4757e-04
1.8162 1.6708e+06 3.7132e-02 1.9390e-02 7.1185e-03
1.7966 1.6776e+06 4.4358e-02 3.7320e-02 2.1884e-02
1.7770 1.6718e+06 4.4311e-02 3.7523e-02 2.0657e-02
1.7639 1.6711e+06 3.3480e-02 1.3616e-02 3.3972e-03
1.7509 1.6727e+06 1.2058e-02 1.2782e-03 4.1071e-04
1.7313 1.7032e+06 1.2193e-02 9.8461e-04 4.0629e-04
1.7118 1.7030e+06 2.7104e-01 3.8657e-02 7.5183e-03
1.6923 1.6879e+06 5.9172e-01 4.8241e-01 2.6812e-01
1.6793 1.6945e+06 6.0731e-01 5.4837e-01 3.5752e-01
1.6532 1.6892e+06 5.1345e-01 2.8133e-01 8.9126e-02
The pressure was set 1.79 bar.
The beam intensity at S1 was set to 1M. |
|
|
|
|
|
|